Suppose, there is a consumer who derives utility from the consumption of two goods, X & Y, her utility function is U (X,Y) = X0.75 Y0.25. Initially, the price of X & Y are given byP_x&P_y. Her total income is given by m.

a) Form the Lagrangian Function for the the expenditure minimization problem.
b) Derive the Hicksian demand functions for X & Y.
c) If U=20, P_x=6 &P_y=2, find the values of X & Y.
d) Based on your answers in part (c) what can you conclude about the relation between the two goods?
1
Expert's answer
2020-10-15T11:58:50-0400

"U(x.y) = X" 0.75 "Y" 0.25


A) "h = min P1X +P1 Y - \lambda(X" 0.75"Y" 0.25 "- \bar U" ")"

"\frac{dh}{dx} = P1-" "\lambda 0.75 X" -0.25"Y" 0.25 ----------- 1


"\frac{dh}{dx} = P2-" "\lambda 0.25 X" -0.75"Y" 0.75 "= 0" ----------- 2


"\frac{dx}{d\lambda} = - (X" 0.75 "Y" 0.25"-" "\bar U)" "= 0" --------------- 3

using equation 1 and 2


"\frac{P1}{P2} =\frac{3Y}{X}" --------------------------------------------------- 4


B)- Put equation 4 and 3


"X"c "= \bar U(\frac{3P2}{P1})" 0.25 , "Y"C "= \bar U(\frac{P1}{3P2})"0.75


C) Putting "\bar U = 20 ,P1 = 6 , P2 = 2" in "X"C and "Y" C

We get "X" C "= 20" ,"Y" C "= 20"


D) part C shows that for any price level of X and Y ,The Hickson demand for X and Y would be same